OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b^2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x^2 + 7x + 3 = 0 Help please! Best answer and a fan!

OpenStudy (anonymous):

@satellite73

hero (hero):

Hint: The part of the quadratic formula that can be used to determine whether or not a quadratic equation can be solved by factoring is also known as the discriminant.

OpenStudy (anonymous):

I've been stuck on this for like 2 days, and I haven't been able to submit my quiz, I don't even know where to start.

OpenStudy (anonymous):

Is it going to be one of these −b b^2 − 4ac 2a or is it going to be a numeral?

OpenStudy (anonymous):

if \(b^2-4ac\) is a perfect square, like say \(25\) or \(49\) then \(\sqrt{b^2-4ac}\) is a whole number, so you have rationals answers, so you can factor

OpenStudy (anonymous):

Is the answer 25?

OpenStudy (anonymous):

if \(b^2-4ac\) is not a perfect square, then you cannot factor using whole numbers or fractions

OpenStudy (anonymous):

no, 25 is not the answer, it is an example of a perfect square

OpenStudy (anonymous):

Ugh, I hate this.

OpenStudy (anonymous):

it is really not such a big deal if the discriminant \(b^2-4ac\) is a square, then you can factor if it is not a square, then you cannot

OpenStudy (anonymous):

the quadratic formula \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] has a square root in it if the number inside the radical is a perfect square, then you will have a fraction if you have a fraction, you can factor

OpenStudy (anonymous):

the answer is 25 Actually.